Main INDEX
Monthly INDEX
PREV
NEXT
Make New Entry,
Make Followup Entry
User name V. Sulkosky
Log entry time 12:47:43 on February 17, 2009
Entry number 261625
Followups:
keyword=Experiment Summary for Feb 16, Owl - Feb 17, Owl
Summary for Feb 16 owl - Feb 17 owl
Smooth data taking until 22:30, then beam instability problems.
Beam in Use
shift |
ABU |
BANU |
BNA |
2009/02/16 owl | 7.3 |
0.17 | 0.53 |
2009/02/16 day | 7.20 |
0.39 | 0.42 |
2009/02/16 swing | 5.84 |
0.64 | 1.52 |
2009/02/17 owl | 0.0 |
0.0 | 8.0 |
total | 20.3 |
1.2 | 10.5 |
Main events/issues:
- Transverse Production
- Feb. 16, beam position and profile check.
- Feb. 16, 01:12-05:30 Complete kinematic M: P_lhrs = 1.7 GeV/c, 45
degrees.
- Left HRS momentum change to 1.6 GeV/c.
- Feb. 16, 05:45 Begin kinematic L: P_lhrs = 1.6 GeV/c, 45 degrees.
- Feb. 16, 17:40 Change target polarization direction to 270 degrees.
- Feb. 16, 22:35 beam instability begins due to Moller quad power
supply.
- Feb. 16 22:35 to Feb. 17 08:00 Trouble shooting and fix of beam
instability problem.
- Total accummulated charge for kinematic L: 0.76 C out of 1.3 C.
Plans for the next day:
- Beam studies: Monday 8:00-14:00
- Controlled Access until 14:00
- Target polarimetry tests
- Switch to longitudinal polarization
- BCM Calibration
- Nitrogen Reference Cell Run at 135 psig
- Begin longitudinal production 11 hours beam on target
- Insert beam half-wave plate for half of the statistics
A copy of this log entry has been emailed to: brads,jpchen,jiang,meziani,choi